OpenStudy (anonymous):

You want to display the hexadecimal number: \[ABC_{16}\] in decimal form. What hexademical number should you add to this so that it will display in hexadecimal the same as if it was converted to decimal? For example, \[A_{16} = 10_{10} and A_{16} + 6_{16} = 10_{16},\] so to make "A" display the same as decimal, you would add \[6_{16}\]

OpenStudy (asnaseer):

first step is to convert \(ABC_{16}\) into decimal. what do you get for this?

OpenStudy (anonymous):

Um, I'm not used to hexadecimals at all. I'm only used to decimals and this is a new topic to me.

OpenStudy (asnaseer):

I suggest you first study this: http://www.mathsisfun.com/hexadecimals.html and then come back to this question.

OpenStudy (anonymous):

From a quick google search though, I think that it's 2748

OpenStudy (anonymous):

Did I get it right?

OpenStudy (asnaseer):

it better to learn the concept than to reply on google. I would really advise you to first study the link I gave you and then come back, otherwise you will just get stuck in the next steps.

OpenStudy (asnaseer):

*to rely on google

OpenStudy (anonymous):

Ok. I learned how to convert hexadecimal to decimal and the columns and counting system before, but do I need to know "the point"?

OpenStudy (phi):

the positions in hex are powers of 16 \[ 16^2 16^1 16^0 \] for the first 3 digits. to accomodate numbers 10 through 15 we use A,B,C,D,E,F A 10 B 11 C12 D13 E14 F15 so ABC in hex is \[A*16^2 + B*16^1 + C*16^0 \] or 10*256 + 11*16 + 12*1 = 2748

OpenStudy (anonymous):

Yeah. I understand that per. But in the example, it says: \[A_{16} + 6_{16} = 10_{16}\] which I can't seem to figure out why it works.

OpenStudy (phi):

To find what to add to ABC hex to get 2748 hex, we subtract (in hex!) \[ 2748_{16} - ABC_{16} \] an interesting exercise, but you borrow 16 rather than 10...

OpenStudy (phi):

\[ A_{16} + 6_{16} = 16_{10} \] but 16 is the next power of 16, so in hex, \[ 10_{16} \]

OpenStudy (anonymous):

Oh ok.

OpenStudy (phi):

so F+F = (15+15 decimal=30 decimal= 1*16+14 --> 1E hex

OpenStudy (phi):

of course, you could memorize the addition table for hex, or count on your 16 fingers(!)

OpenStudy (anonymous):

2748 -ABC ----- 1C8C I that it?

OpenStudy (phi):

that's what I got

OpenStudy (anonymous):

Alright. That's what the answer key says too. Thanks! It was pretty simple once I switched the system of 10 to 16 :)